How much calcium chloride is needed to make 1.00 liter of a 1 M solution of calcium chloride?

A solution contains 33.2 g H3PO4 in 440. mL of solution. Calculate its molarity.

What is the molarity of a hydrochloric acid solution that contains 18.4 g HCl in 102.2 mL of solution?
1
Expert's answer
2020-05-10T09:19:56-0400

Molecular mass of CaCl2 equal:

M=111g/mol

There are 111g CaCl2 in 1L solution with molarity 1M


n(H3PO4)=33,2g/98g/mol=0,34mol

с=0,34mol/0,44L=0,77M


n(HCl)=18,4g/36,46g/mol=0,5mol

c=0,5/0,1022L=4,89M